OpenStudy (anonymous):

Hi I need your help with this! What's the Cartesian equation for this: r = (3)+(-5)*COS(θ)

OpenStudy (amistre64):

x = r cos(theta) .... might be useful

OpenStudy (amistre64):

r = (3)+(-5)*COS(θ) multiply by r r^2 = 3r -5 rCOS(θ) r^2 = 3r - 5x ; r^2 = x^2 + y^2 x^2 + y^2 = 3r - 5x; any way to get rid of that pesky r?

OpenStudy (amistre64):

x^2 + y^2 = 3sqrt(x^2+y^2) - 5x x^2 + 5x + y^2 = 3sqrt(x^2+y^2) (x^2 + 5x + y^2)^2 = 9x^2 + 9y^2 ... maybe

OpenStudy (anonymous):

Hey! I do not mean to sound ungrateful for not replying so I had to say this. THANK YOU. THANK YOU SO MUCH. Honestly you saved my life right there, I was so lost but when I read you answer I understood and thank you so much for the visual on the image. I thought it would look like that but I was unsure. You are really awesome amistre64! Again thank you.